OpenStudy (anonymous):

The length of a rectangle is 5.5 feet. If the area is 44 square feet, what is the width of the rectangle, in feet?

OpenStudy (sweetburger):

8ft

OpenStudy (sweetburger):

44/5.5 =8 because l x w= A so 5.5 x w =44 divide 5.5 and that is your width

OpenStudy (cwrw238):

Area = width * length plugging in the values : 44 = w * 5.5 where w = width divide both sides by 5.5: 44/5.5 = w w = 8 ft ;